OpenStudy (anonymous):

A right circular cylinder has a base whose area is 616 square feet. If the height of the cylinder is 9 feet, what is its lateral surface area? Use π

OpenStudy (anonymous):

because the base whose area is 616 square feet,which Sb=pi*r^2 r=sqrt{S/(pi)}=sqrt{616/pi} lateral surface area=perimeter of the base* height=2pi*r*h=2pi*sqrt{616/pi}*9 put this into the caculator finished!

OpenStudy (anonymous):

Idk if you calculated, but I got 1,408. is that correct?

OpenStudy (anonymous):

i don't have one nearby,but i think you are wrong, i guess it's around 735

OpenStudy (anonymous):

hmmm, I recalculated and got 795. Does that seem right?

OpenStudy (anonymous):

yep,if pi 3.14, the answer is 791. your answer is right!

OpenStudy (anonymous):

Thank you so much! :) You're awesome.

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

Wouldn't it be 792?